OpenStudy (anonymous):

please and thank you! Choose the correct simplification of the expression c2 • c9.

OpenStudy (diyadiya):

Use this \[\large x^m \times x^n = x^{m+n}\]

OpenStudy (diyadiya):

Your previous question was very similar to this one

OpenStudy (anonymous):

is it c^11?

OpenStudy (diyadiya):

Right! :)

OpenStudy (anonymous):

(a2)3 and this is a^5?

OpenStudy (diyadiya):

No \[\large ((x)^m)^n= x^{m \times n}\]

OpenStudy (apoorvk):

no if you mean a^2 times 3. also no if you mean (a^2)^3

OpenStudy (anonymous):

ugggg i dont get (a2)3............. :(

OpenStudy (diyadiya):

\[\large ((a)^2)^3= a^{2 \times 3 }\]

OpenStudy (anonymous):

when do u get out of school @gurrilabill

OpenStudy (anonymous):

june 10th @drsuz98

OpenStudy (anonymous):

woah! My last day was today but Im not done..i have until moday! @gurrilabill